OpenStudy (anonymous):

Find dy/dx with implicit differentiation (xy)^1/2=1+x^2y

OpenStudy (ranga):

use power rule, chain rule and product rule.

OpenStudy (anonymous):

Can you show me?

OpenStudy (amistre64):

let a = xy a' = x'y + xy' by product rule let b = x^2y ... which is either menas x^2 y or x^(2y) hard to tell b' = derive it as needed then assess: a^(1/2) = 1 + b 1/2 a^(-1/2) a' = 1 + b'

OpenStudy (anonymous):

It's x^2 y

OpenStudy (amistre64):

then thats pretty much a product rule as well ... how would you define b' with that?

OpenStudy (anonymous):

2xy+x^2 dy/dx ?

OpenStudy (amistre64):

correct .. but i like the ' notations b' = 2x x' y + x^2 y' the reason i keep the x' parts is to show the chain rule in action. now x' = dx/dx = 1 to clean it up afterwards

OpenStudy (amistre64):

let a = xy a' = x'y + xy' by product rule let b = x^2 y b' = 2x x' y + x^2 y' a^(1/2) = 1 + b 1/2 a^(-1/2) a' = 1 + b' the ab was just for a cleanup ... to easier see an overall picture, sub back in the information 1/2 (xy)^(-1/2) (y+xy') = 1 + (2xy + x^2 y') the rest is just algebra to solve for y'

OpenStudy (anonymous):

Okay I get it now thank you very much :)